### v2

# Selecting the Number of Communities for Weighted Degree-Corrected Stochastic Block Models

Abstract
We investigate how to select the number of communities for weighted networks without a full likelihood modeling. First, we propose a novel weighted degree-corrected stochastic block model (DCSBM), in which the mean adjacency matrix is modeled as the same as in standard DCSBM, while the variance profile matrix is assumed to be related to the mean adjacency matrix through a given variance function. Our method of selecting the number of communities is based on a sequential testing framework, and in each step the weighted DCSBM is fitted via some spectral clustering method. A key step is to carry out matrix scaling on the estimated variance profile matrix. The resulting scaling factors can be used to normalize the adjacency matrix, from which the testing statistic is obtained. Under mild conditions on the weighted DCSBM, our proposed procedure is shown to be consistent in estimating the true number of communities. Numerical experiments on both simulated and real-world network data also demonstrate the desirable empirical properties of our method.
Keywords: rank selection, weighted networks, DCSBM, sequential testing, variance profile, matrix scaling
1 Introduction
In network analysis, community detection is regarding how to partition the vertices of a graph into clusters with similar connection patterns. This problem has numerous applications in a variety of fields, including applied physics, sociology, economics, and biology; see the survey paper Fortunato [2010]. An important problem of community detection is how to select the number of communities. To answer this question, various methods have been proposed for unweighted networks, such as sequential testing [Lei, 2016; Jin et al., 2022; Han et al., 2023], spectral thresholding [Le and Levina, 2015], and model comparison based methods [Daudin et al., 2008; Wang and Bickel, 2017; Saldana et al., 2017; Chen and Lei, 2018; Li et al., 2020; Hu et al., 2020; Ma et al., 2021], etc. In particular, sequential testing methods are usually based on goodness-of-fit testing methods [e.g., Gao and Lafferty, 2017; Jin et al., 2018, 2021; Hu et al., 2021].
Most existing methods of selecting the number of communities are only applicable to unweighted networks. In contrast, selecting the number of communities for weighted networks is little studied. In fact, weighted networks are common in practice, and usually reveal refined community structures; see, for example, Newman and Girvan [2004], Ball et al. [2011] and Newman and Reinert [2016]. A challenge for rank selection in weighted networks is on the modeling of the networks, since likelihood based methods could be very restrictive for this problem.
[…]
However, such explicit and tight bound $\|\bm{A}-\bm{M}\|\leq\tau$ is in general not available. One natural idea is to consider a normalized version of the noise matrix, such that a tight and explicit upper bound for its spectral norm can be derived. Here we borrow an idea from Landa et al. [2022], which studies how to reveal the rank of a Poisson data matrix by spectral truncation with the assistance of matrix scaling on the variance profile matrix. Coming back to our problem, if the variance profile matrix $\bm{V}$ is known, whose entries are $V_{ij}=\textrm{var}(A_{ij})=\textrm{var}(A_{ij}-M_{ij})$. It is known that there exists a diagonal matrix $\bm{\Psi}=\operatorname{Diag}(\psi_{1},\ldots,\psi_{n})$, such that $\bm{\Psi}\bm{V}\bm{\Psi}$ is doubly stochastic, i.e., every row sum of $\bm{\Psi}\bm{V}\bm{\Psi}$ is $1$ [e.g., Knight et al., 2014]. With this scaling diagonal matrix $\bm{\Psi}$, recent results [e.g., Latała et al., 2018] in random matrix theory guarantees $\left\|\bm{\Psi}^{\frac{1}{2}}(\bm{A}-\bm{M})\bm{\Psi}^{\frac{1}{2}}\right\|\leq 2+\epsilon$ for some small positive constant $\epsilon$.
Based on the above heuristic, we propose the following sequential testing procedure: for each $m=1,2,3,\ldots$, in the $m$-th step of the procedure, we first group the nodes into $m$ distinct communities via some spectral clustering method, e.g., SCORE proposed in Jin [2015] or regularized spectral clustering (RSC) [Amini et al., 2013; Joseph and Yu, 2016]. With the estimated groups, we can obtain the estimated mean adjacency matrix $\widehat{\bm{M}}^{(m)}$ by fitting the DCSBM, and further obtain the estimated variance profile matrix $\widehat{\bm{V}}^{(m)}$ by using the variance-mean relationship. Next, we find a diagonal scaling matrix $\widehat{\bm{\Psi}}^{(m)}$ such that $\widehat{\bm{\Psi}}^{(m)}\widehat{\bm{V}}^{(m)}\widehat{\bm{\Psi}}^{(m)}$ is doubly stochastic. The test statistic is chosen as $T_{n,m}=\left|\lambda_{m+1}\left(\left(\widehat{\bm{\Psi}}^{(m)}\right)^{\frac{1}{2}}\bm{A}\left(\widehat{\bm{\Psi}}^{(m)}\right)^{\frac{1}{2}}\right)\right|$. For $m=1,2,3,\ldots$, we stop the iterative procedure if $T_{n,m}<2+\epsilon$, and then report the estimated number of communities as $\widehat{K}=m$.
1.1 Related Work
Stochastic block models (SBM) [Holland et al., 1983] and its variants such as degree-corrected stochastic block models (DCSBM) [Karrer and Newman, 2011] have been widely used to model community structures in networks. Many community detection methods have been proposed in the literature, with examples including modularity and likelihood based methods [e.g., Newman and Girvan, 2004; Bickel and Chen, 2009; Zhao et al., 2012; Amini et al., 2013; Bickel et al., 2013; Le et al., 2016; Zhang and Zhou, 2016, 2020], spectral methods [e.g., Rohe et al., 2011; Amini et al., 2013; Jin, 2015; Lei and Rinaldo, 2015; Joseph and Yu, 2016; Gulikers et al., 2017; Gao et al., 2018; Abbe et al., 2020; Zhang et al., 2020; Deng et al., 2021; Li et al., 2022; Lei et al., 2020], convex optimization methods [e.g., Cai and Li, 2015; Guédon and Vershynin, 2015; Abbe et al., 2015; Chen et al., 2018; Li et al., 2021], and many other methods such as Bayesian approaches.
Matrix scaling has been used in Landa et al. [2022] to reveal the rank of a Poisson data matrix, where the variance profile matrix is the same as the mean adjacency matrix, and thereby can be estimated by the observed count data matrix. In contrast, our work is not restricted to Poisson data matrix. On the other hand, we have a parametric model for the mean structure in our weighted DCSBM, while there is no such parametric mean modeling in Landa et al. [2022].
Sequential testing based on stepwise SBM/DCSBM fitting has been commonly used in the literature of rank selection for standard binary networks; see, e.g., Lei [2016]; Wang and Bickel [2017]; Hu et al. [2020]; Ma et al. [2021]; Hu et al. [2021]; Jin et al. [2022]. However, these methods either rely on the binary structure of unweighted networks, or on modeling the complete likelihood of the random network, and thus cannot be straightforwardly extended to our setup.
To address some technical challenges for the underfitting case $m<K$, we generalize the Nonsplitting Property established in Jin et al. [2022] for spectral clustering approach SCORE proposed in Jin [2015] to weighted DCSBM. This is the major reason why our main results are established when the spectral method is chosen as SCORE, though extensions to other spectral methods may be possible, which is left for future investigation.
1.2 Organization of the Paper
Our paper is organized as follows. In Section 2, we extend the standard DCSBM to model the mean structure of weighted networks, and link the mean adjacency matrix and the variance profile matrix with a variance function. In Section 3, we introduce a stepwise DCSBM fitting procedure, with a spectral test statistic based on variance profile matrix scaling. Our main results on the consistency of our procedure are stated in Section 4 under certain assumptions on the weighted DCSBM. In Section 5, we demonstrate the empirical properties of our proposed procedure with both synthetic and real-world weighted networks. The proofs of our main results are given in Appendix B, with some preliminaries and proofs of technical lemmas also included in the rest of the appendix.
2 Weighted Degree-Corrected Stochastic Block Models
Our first task is to extend the standard DCSBM proposed in Karrer and Newman [2011] to networks with nonnegative weights. Rather than specifying the entire likelihood of the random graph with weighted edges, our weighted DCSBM relies on specifying the first two moments, i.e., the mean adjacency matrix and the variance profile matrix.
We first introduce some notations for the community structure. Let $\bm{A}$ be the adjacency matrix of the weighted network with $n$ nodes, which belong to $K$ separate communities. Denote by $\mathcal{N}_{1},\ldots,\mathcal{N}_{K}$ the underlying communities, with respective cardinalities $n_{1},\ldots,n_{K}$, and thereby $n=n_{1}+\cdots+n_{K}$. Denote by $\phi:[n]\rightarrow[K]$ the community membership function of nodes, such that $\phi(i)=k$ if and only if node $i$ belongs to community $\mathcal{N}_{k}$. We also use vector $\bm{\pi}_{i}\in\mathbb{R}^{K}$ to indicate the community belonging of node $i$ by letting $\pi_{i}(k)=1$ if $i\in\mathcal{N}_{k}$ and $\pi_{i}(k)=0$ otherwise. Denote $\bm{\Pi}=[\bm{\pi}_{1},\ldots,\bm{\pi}_{n}]^{\top}\in\mathbb{R}^{n\times K}$ as the community membership matrix.
The mean adjacency matrix is modeled in a similar manner as in the standard DCSBM [Karrer and Newman, 2011]. Note that the network is allowed to have self-loops for simplicity of analysis without loss of generality. Concretely, let $\bm{B}$ be a $K\times K$ symmetric community connectivity matrix with positive entries and diagonal entries $B_{kk}=1$ for $k=1,\ldots,K$ to ensure identifiability. Also, let $\theta_{1},\ldots,\theta_{n}>0$ be the heterogeneity parameters. Denote $\bm{\theta}=(\theta_{1},\ldots,\theta_{n})^{\top}$ and $\bm{\Theta}=\text{diag}(\theta_{1},\ldots,\theta_{n})$. Then, the entries of the expected adjacency matrix are parameterized as
$\operatorname{\mathbb{E}}[A_{ij}]\coloneqq M_{ij}\coloneqq\theta_{i}\theta_{j}B_{\phi(i)\phi(j)}=\theta_{i}\theta_{j}\bm{\pi}_{i}^{\top}\bm{B}\bm{\pi}_{j},\quad 1\leq i\leq j\leq n.$ (1)
In matrix form, the expected adjacency matrix can be represented as
$\bm{M}=\bm{\Theta}\bm{\Pi}\bm{B}\bm{\Pi}^{\top}\bm{\Theta}.$ (2)
The weighted edges are assumed to be independent. In modeling the second order moments, we denote $V_{ij}=\textrm{var}(A_{ij})$, and assume that it is a function of the corresponding mean by some variance function $\nu(\cdot)$, i.e., $V_{ij}=\nu(M_{ij})$. By denoting the variance profile matrix $\bm{V}=[V_{ij}]_{i,j=1}^{n}$, we can also represent the variance-mean relationship as $\bm{V}=\nu(\bm{M})$, where $\nu$ is viewed as an entrywise operation.
The variance function $\nu(\cdot)$ is known for some common distributions, such as $\nu(\mu)=\mu(1-\mu)$ for the Bernoulli case, and $\nu(\mu)=\mu$ for the Poisson case. We assume $\nu$ is known in the present work. It would be interesting to study how to estimate $\nu$ in a data-dependent manner, but we leave this for future investigation.
3 Methodology
In this section, we will provide the details of our stepwise procedure for estimating the number of communities in weighted networks. Our method of estimating the number of communities is a stepwise method. It relies on fitting the weighted DSCBM with an increasing sequence of candidate number of communities. In other words, for each positive integer $m=1,2,\ldots$, we aim to test $H_{0}:K=m$, where $K$ is the true rank of the weighted DCSBM.
3.1 Stepwise Weighted DCSBM Fitting
With each hypothetical number of communities $m$, we apply some standard community detection method, such as SCORE [Jin, 2015] or regularized spectral clustering (RSC) [Amini et al., 2013; Joseph and Yu, 2016], to obtain $m$ estimated communities $\widehat{\mathcal{N}}_{1}^{(m)},\ldots,\widehat{\mathcal{N}}_{m}^{(m)}$. Then, the DCSBM parameters $\bm{\theta}$, $\bm{B}$ and $\bm{W}$ can be directly estimated in a plug-in manner.
The estimates of $\bm{\theta}$ and $\bm{B}$ are the same as in the case of standard DCSBM, and we give a review of their derivations in Jin et al. [2022]. Let’s first see how to represent $\bm{\theta}$ and $\bm{B}$ with the mean adjacency matrix $\bm{M}$, the true communities $\mathcal{N}_{1},\ldots,\mathcal{N}_{K}$, and the expected degrees. Decompose the vector $\bm{\theta}$ as $\bm{\theta}=\bm{\theta}_{1}+\cdots+\bm{\theta}_{K}$, where $\bm{\theta}_{k}\in\mathbb{R}^{n}$ for $k=1,\ldots,K$ such that $\bm{\theta}_{k}(i)=\theta_{i}$ if $i\in\mathcal{N}_{k}$ and $\bm{\theta}_{k}(i)=0$ otherwise. We can similarly decompose the $n$-dimensional all-one vector into $\bm{1}_{n}=\bm{1}_{1}+\cdots+\bm{1}_{K}$ such that $\bm{1}_{k}(j)=1$ if $j\in\mathcal{N}_{k}$ and $\bm{1}_{k}(j)=0$ otherwise. It is easy to verify that for $1\leq k,l\leq K$, $\bm{1}_{k}^{\top}\bm{M}\bm{1}_{l}=B_{kl}\|\bm{\theta}_{k}\|_{1}\|\bm{\theta}_{l}\|_{1}$. By the assumption $B_{kk}=1$ for $k=1,\ldots,K$, the above equality implies $\|\bm{\theta}_{k}\|_{1}=\sqrt{\bm{1}_{k}^{\top}\bm{M}\bm{1}_{k}}$, which further gives
$B_{kl}=\frac{\bm{1}_{k}^{\top}\bm{M}\bm{1}_{l}}{\|\bm{\theta}_{k}\|_{1}\|\bm{\theta}_{l}\|_{1}}=\frac{\bm{1}_{k}^{\top}\bm{M}\bm{1}_{l}}{\sqrt{\bm{1}_{k}^{\top}\bm{M}\bm{1}_{k}}\sqrt{\bm{1}_{l}^{\top}\bm{M}\bm{1}_{l}}}.$ (3)
Denote the degree corresponding of node $i\in\mathcal{N}_{k}$ as $d_{i}=\sum_{j=1}^{n}A_{ij}$, whose expectation, referred to as the population degree, is
$d_{i}^{*}\coloneqq\operatorname{\mathbb{E}}[d_{i}]$ $=\theta_{i}\left(B_{k1}\|\bm{\theta}_{1}\|_{1}+B_{k2}\|\bm{\theta}_{2}\|_{1}+\cdots+B_{kK}\|\bm{\theta}_{K}\|_{1}\right)$
$=\theta_{i}\left(\frac{\bm{1}_{k}^{\top}\bm{M}\bm{1}_{1}}{\sqrt{\bm{1}_{k}^{\top}\bm{M}\bm{1}_{k}}}+\cdots+\frac{\bm{1}_{k}^{\top}\bm{M}\bm{1}_{K}}{\sqrt{\bm{1}_{k}^{\top}\bm{M}\bm{1}_{k}}}\right)=\theta_{i}\frac{\bm{1}_{k}^{\top}\bm{M}\bm{1}_{n}}{\sqrt{\bm{1}_{k}^{\top}\bm{M}\bm{1}_{k}}}.$
This implies that the degree-correction parameter $\theta_{i}$ can be expressed as
$\theta_{i}=\frac{\sqrt{\bm{1}_{k}^{\top}\bm{M}\bm{1}_{k}}}{\bm{1}_{k}^{\top}\bm{M}\bm{1}_{n}}d_{i}^{*},\quad i\in\mathcal{N}_{k}.$ (4)
With (3) and (4), we obtain plug-in estimates of $\bm{\theta}$ and $\bm{B}$ by replacing the true community partition $\mathcal{N}_{1},\ldots,\mathcal{N}_{K}$ with the estimated partition $\widehat{\mathcal{N}}_{1}^{(m)},\ldots,\widehat{\mathcal{N}}_{m}^{(m)}$, replacing $\bm{M}$ with $\bm{A}$, and replacing $d_{i}^{*}$ with $d_{i}$. In analogy to $\bm{1}_{n}=\bm{1}_{1}+\cdots+\bm{1}_{K}$, we decompose the all-one vector as the sum of indicator vectors corresponding to the estimated communities: $\bm{1}_{n}=\hat{\bm{1}}_{1}^{(m)}+\cdots+\hat{\bm{1}}_{m}^{(m)}$, where for each $j=1,\ldots,n$ and $k=1,\ldots,m$, $\hat{\bm{1}}_{k}^{(m)}(j)=1$ if $j\in\widehat{\mathcal{N}}_{k}^{(m)}$ and $\hat{\bm{1}}_{k}^{(m)}(j)=0$ otherwise. Then, the plug-in estimates are
$\hat{\theta}^{(m)}_{i}\coloneqq\frac{\sqrt{(\hat{\bm{1}}_{k}^{(m)})^{\top}\bm{A}\hat{\bm{1}}_{k}^{(m)}}}{(\hat{\bm{1}}_{k}^{(m)})^{\top}\bm{A}\bm{1}_{n}}d_{i},\quad k=1,\ldots,m\text{~{}and~{}}i\in\widehat{\mathcal{N}}_{k}^{(m)},$ (5)
and
$\widehat{B}_{kl}^{(m)}\coloneqq\frac{(\hat{\bm{1}}^{(m)}_{k})^{\top}\bm{A}\hat{\bm{1}}^{(m)}_{l}}{{\|\hat{\bm{\theta}}^{(m)}_{k}\|_{1}\|\hat{\bm{\theta}}^{(m)}_{l}\|_{1}}}=\frac{(\hat{\bm{1}}^{(m)}_{k})^{\top}\bm{A}\hat{\bm{1}}^{(m)}_{l}}{\sqrt{(\hat{\bm{1}}^{(m)}_{k})^{\top}\bm{A}\hat{\bm{1}}^{(m)}_{k}}\sqrt{(\hat{\bm{1}}^{(m)}_{l})^{\top}\bm{A}\hat{\bm{1}}^{(m)}_{l}}},\quad 1\leq k,l\leq m.$ (6)
Then we have the estimated mean matrix: For any $1\leq i,j\leq n$ with $i\in\widehat{\mathcal{N}}_{k}^{(m)}$ and $j\in\widehat{\mathcal{N}}_{l}^{(m)}$,
$\widehat{M}_{ij}^{(m)}\coloneqq\hat{\theta}_{i}^{(m)}\hat{\theta}_{j}^{(m)}\widehat{B}_{kl}^{(m)}=\frac{(\hat{\bm{1}}^{(m)}_{k})^{\top}\bm{A}\hat{\bm{1}}^{(m)}_{l}}{(\hat{\bm{1}}_{k}^{(m)})^{\top}\bm{A}\bm{1}_{n}\cdot(\hat{\bm{1}}_{l}^{(m)})^{\top}\bm{A}\bm{1}_{n}}d_{i}d_{j}.$ (7)
If we denote the community membership matrix corresponding to $\widehat{\mathcal{N}}_{1}^{(m)},\ldots,\widehat{\mathcal{N}}_{m}^{(m)}$ as $\widehat{\bm{\Pi}}^{(m)}\in\mathbb{R}^{n\times m}$, the matrix form of mean estimation is (omitting the superscripts)
$\widehat{\bm{M}}^{(m)}=\widehat{\bm{\Theta}}\widehat{\bm{\Pi}}\widehat{\bm{B}}\widehat{\bm{\Pi}}^{\top}\widehat{\bm{\Theta}}.$ (8)
With the estimated mean matrix $\widehat{\bm{M}}^{(m)}$, we have the residual matrix $\bm{R}^{(m)}=\bm{A}-\widehat{\bm{M}}^{(m)}$ and the estimated variance profile matrix $\widehat{\bm{V}}^{(m)}=\nu(\widehat{\bm{M}}^{(m)})$. In the case where the variance function is unknown, we replace $\nu(\cdot)$ with the estimated variance function $\widehat{\nu}(\cdot)$.
3.2 Variance Profile Scaling and Spectral Statistic
We have introduced in Section 1 how to define the spectral test statistic based on variance profile matrix scaling. We briefly introduce it again in this section. For the variance profile matrix $\bm{V}$, assume $\bm{\Psi}=\operatorname{Diag}(\bm{\psi})=\operatorname{Diag}(\psi_{1},\ldots,\psi_{n})$ is a diagonal matrix such that $\bm{\Psi}\bm{V}\bm{\Psi}$ is doubly stochastic, i.e.,
$\sum_{i=1}^{n}V_{ij}\psi_{i}\psi_{j}=1,\quad\forall j=1,\ldots,n.$ (9)
Uniqueness and existence of such $\bm{\Psi}$ and related algorithms are well-known in the literature; see, e.g., Sinkhorn [1967]; Knight et al. [2014].
Analogously, in the $m$-step of our procedure, we choose scaling factors $\hat{\psi}_{1}^{(m)},\ldots,\hat{\psi}_{n}^{(m)}$, such that $\widehat{\bm{\Psi}}^{(m)}\widehat{\bm{V}}^{(m)}\widehat{\bm{\Psi}}^{(m)}$ is doubly stochastic, where $\widehat{\bm{\Psi}}^{(m)}=\operatorname{Diag}(\hat{\psi}_{1}^{(m)},\ldots,\hat{\psi}_{n}^{(m)})$. In other words,
$\sum_{i=1}^{n}\widehat{V}_{ij}\hat{\psi}_{i}\hat{\psi}_{j}=1,\quad\forall j=1,\ldots,n.$ (10)
Define the test statistic as
$T_{n,m}=\left|\lambda_{m+1}\left(\left(\widehat{\bm{\Psi}}^{(m)}\right)^{\frac{1}{2}}\bm{A}\left(\widehat{\bm{\Psi}}^{(m)}\right)^{\frac{1}{2}}\right)\right|.$
For $m=1,2,3,\ldots$, we stop the iterative procedure if $T_{n,m}<2+\epsilon$ for some prespecified small constant $\epsilon$, and then report the estimated number of communities as $\widehat{K}=m$.
4 Main Results
We aim to establish in this section the consistency of our proposed procedure for selecting the number of communities under the weighted DCSBM. Our main results consist of two parts: (1) In the null case $m=K$, we aim to show that $T_{n,m}\leq 2+o_{p}(1)$; (2) In the underfitting case $m<K$, we will show that $T_{n,m}\gg O_{p}(1)$. We first introduce a sequence of conditions on the weighted DCSBM under which our consistency analysis is conducted.
4.1 Assumptions
The conditions imposed on the weighted DCSBM are captured by a constant $c_{0}$.
Assumption 1.
We consider the weighted DCSBM described in Section 2. Denote $\theta_{\max}=\max\{\theta_{1},\ldots,\theta_{n}\}$ and $\theta_{\min}=\min\{\theta_{1},\ldots,\theta_{n}\}$. Assume the following conditions hold:
  • [Fixed rank] The true number of communities $K$ is assumed to be fixed.
  • [Balancedness]
    $\min_{1\leq k\leq K}\frac{n_{k}}{n}\geq c_{0}\quad\text{and}\quad\frac{\theta_{\min}}{\theta_{\max}}\geq c_{0}.$ (11)
  • [Sparseness]
    $\frac{1}{c_{0}}\geq\theta_{\max}\geq\theta_{\min}\geq\frac{\log^{3}n}{\sqrt{n}}.$ (12)
  • [Community connectivity] The $K\times K$ matrix $\bm{B}$ is fixed, and its entries and eigenvalues are assumed to satisfy
    $\begin{cases}B_{kk}=1\text{~{}~{}~{}for~{}~{}}k=1,\ldots,K,\quad\text{and~{}}c_{0}\leq B_{kl}\leq 1\text{~{}~{}for~{}}1\leq k,l\leq K,\\ \lambda_{1}(\bm{B})>|\lambda_{2}(\bm{B})|\geq\cdots\geq|\lambda_{K}(\bm{B})|\geq c_{0}>0.\end{cases}$ (13)
  • [Variance-mean function] The function $\nu(\cdot)$ satisfies
[…]
  • [Bernstein condition] For any $i\leq j$ and any integer $p\geq 2$, there holds
    $\operatorname{\mathbb{E}}\left[|A_{ij}-M_{ij}|^{p}\right]\leq\left(\frac{p!}{2}\right)R(c_{0})^{p-2}\nu(M_{ij}),$ (15)
    where $R(c_{0})$ is a constant only depending on $c_{0}$.
In fact, the assumption (15) is the standard Bernstein condition, which holds for various common distributions. As an example, the following result shows that the Poisson distribution satisfies this condition.
Proposition 1.
Let $X\sim\text{Poisson}(\lambda)$, where $\lambda\leq C(c_{0})$ and $C(c_{0})$ is a constant only depending on $c_{0}>0$. Then, for any integer $p\geq 2$, there holds
$\operatorname{\mathbb{E}}\left[|X-\lambda|^{p}\right]\leq\left(\frac{p!}{2}\right)R(c_{0})^{p-2}\lambda,$
where $R(c_{0})$ is a constant only depending on $c_{0}$.
Proof.
Since a Poisson random variable is known to be discrete log-concave, by noting that $\textrm{Var}(X)=\lambda$, this lemma can be directly obtained by Lemma 7.5 and Definition 1.2 of Schudy and Sviridenko [2011]. ∎
4.2 Consistency
In our theoretical results introduced in the next section, we choose SCORE [Jin, 2015] for the spectral clustering: First, compute the $m$ leading singular vectors $\bm{u}_{1},\ldots,\bm{u}_{m}$ of $\bm{A}$ corresponding to the $m$ largest eigenvalues in magnitude; next, construct a $n\times(m-1)$ matrix of entrywise ratios $\bm{R}^{(m)}$: $R^{(m)}(i,k)=u_{k+1}(i)/u_{1}(i)$ for $1\leq i\leq n$ and $10\leq k\leq m-$ 1; finally, the rows of the ratio matrix $\bm{R}^{(m)}$ are clustered by the $k$-means algorithm assuming there are $m$ clusters. In future work, we are investigated in establishing the consistency for other commonly used spectral clustering method, e.g., the regularized spectral clustering methods proposed and studied in Amini et al. [2013]; Joseph and Yu [2016].
Now we introduce the main results that guarantee the consistency of the procedure introduced in Section 3 for selecting the number of communities, provided Assumption 1 holds. We separate our main results into two parts: the null case ($m=K$) and the underfitting case ($m<K$).
Theorem 4.1 (Null Case).
If we implement the procedure in Section 3 with candidate number of communities $m=K$ and the spectral clustering chosen as SCORE, then, for any fixed $c_{0}>0$ in Assumption 1, as $n\rightarrow\infty$, we have $T_{n,m}\leq 2+o_{P}(1)$.
Theorem 4.2 (Underfitting Case).
If we implement the procedure in Section 3 with the candidate number of communities $m<K$ and the spectral clustering method chosen as SCORE, then, for any fixed $c_{0}>0$ in Assumption 1, as $n\rightarrow\infty$, we have $T_{n,m}\stackrel{{\scriptstyle P}}{{\longrightarrow}}\infty$.
Obviously, combining Theorems 4.1 and 4.2 shows the consistency of our procedure for estimating the true number of communities $K$. In (11), the sparsity assumption $\theta_{\min}\geq\left.(\log^{3}n)\middle/\sqrt{n}\right.$ might be suboptimal. In contrast, the corresponding assumption is typically $\theta_{\min}\geq C_{0}\sqrt{(\log n)/n}$ in the literature of unweighted network model selection [e.g., Jin et al., 2022]. The extra logarithms for the general DCSBM are majorly due to an application of the truncation technique in the proof, since existing spectral radius results for sparse and heterogeneous random matrices in the literature, e.g., Latała et al. [2018], usually require the entries to be uniformly bounded. It is an interesting question whether the logarithm factors can be further improved, but we leave this for future investigation.
5 Numerical Experiments
We conduct numerical experiments using both synthetic and real-world data to demonstrate the empirical behavior of our proposed method, referred to as Stepwise Variance Profile Scaling (SVPS). In the step of weighted DCSBM fitting in SVPS, we use either SCORE proposed in Jin [2015] or regularized spectral clustering (RSC) proposed in Amini et al. [2013]. If RSC is used, the regularization parameter is set to $0.25\times\bar{d}/n$, where $\bar{d}$ is the average node degree of the network.
Throughout all our experiments, we choose the variance function as $\hat{\nu}(\mu)=\mu$. Note that this does not imply the true variance function satisfies this relationship. In other words, model mismatching is allowed in our experiments. Therefore, the estimated variance profile matrix always satisfies $\widehat{\bm{V}}^{(m)}=\widehat{\bm{M}}^{(m)}=\widehat{\bm{\Theta}}\widehat{\bm{\Pi}}\widehat{\bm{B}}\widehat{\bm{\Pi}}^{\top}\widehat{\bm{\Theta}}$.
The threshold to compare $T_{n,m}$ with in sequential testing is either $2.05$ or $2.10$. For the threshold in sequential testing, we compare $T_{n,m}$ with $2.05$ or $2.10$. We will specify our choices in each specific case of the experiments.
5.1 Synthetic Networks
5.1.1 Weighted DCSBM Generation
Let’s discuss here how to generate the weighted networks based on certain weighted DCSBM for our simulations. We consider three types of distributions: Poisson, binomial, and negative binomial. In each case, the distribution of the weighted DCSBM is determined by the mean structure. Therefore, we first generate the mean matrix following the setup proposed in Hu et al. [2020].
In fact, to determine the mean structure, we only need to specify the $K\times K$ matrix $\bm{B}$, the degree-correction parameters $\bm{\theta}$, and the block sizes. We specify them as follows:
  • For $\bm{B}$, we assume
    $B_{kl}=\rho\left(1+r\times\mathbf{1}(k=l)\right),$
    where $\rho$ is a sparsity parameter and $r$ is a contrast parameter. We consider the following combinations of $(\rho,r)$: $(0.04,4)$, $(0.06,3)$ and $(0.12,2)$ in our simulations.
  • The block sizes are set according to the sequence $(50,100,150,50,100,150)$. For example, $(n_{1},n_{2})=(50,100)$ when $K=2$, $(n_{1},n_{2},n_{3})=(50,100,150)$ when $K=3$, etc.
  • The degree-correction parameters are i.i.d. generated from the following distribution:
    $\begin{cases}\text{Uniform}(0.6,1.4),\quad&\text{with probability}0.8;\\ 0.5,\quad&\text{with probability }0.1;\\ 1.5,\quad&\text{with probability }0.1.\end{cases}$
    After generating $\bm{M}$, we consider the following simulations to generate $\bm{A}$:
    • Simulation 1: Generate $A_{ij}\stackrel{{\scriptstyle ind}}{{\sim}}\text{Poisson}(M_{ij})$ for $1\leq i<j\leq n$.
    • Simulation 2: Generate $A_{ij}\stackrel{{\scriptstyle ind}}{{\sim}}\text{Binomial}(5,M_{ij}/5)$ for $1\leq i<j\leq n$.
    • Simulation 3: Generate $A_{ij}\stackrel{{\scriptstyle ind}}{{\sim}}\text{NB}(5,M_{ij}/5)$ for $1\leq i<j\leq n$, where NB$(n,p)$ is the negative binomial distribution with $n$ trials and success probability $p$ in each trial.
5.1.2 Comparison to Score Based Methods
In our simulated experiments, we compare SVPS with two score based methods: corrected BIC (CBIC) in Hu et al. [2020] and integrated classification likelihood (ICL) method in Daudin et al. [2008], both of which are adapted to weighted DCSBM with explicit likelihoods. To be specific, for each $m=1,2,3,...$, we calculate the following two scores:
$\text{CBIC}(m)=\log f(\bm{A}|\widehat{\bm{M}}^{(m)})-\left[\lambda n\log m+\frac{m(m+1)}{2}\log n\right],$
and
$\text{ICL}(m)=\log f(\bm{A}|\widehat{\bm{M}}^{(m)})-\left[\sum_{k=1}^{m}\hat{n}_{k}\log\left(\frac{n}{\hat{n}_{k}}\right)+\frac{m(m+2)}{2}\log n\right],$
where $\hat{n}_{k}$’s are estimated block sizes according to $\widehat{\bm{M}}^{(m)}$, and the likelihood function is based on Poisson DCSBM, binomial DCSBM or negative binomial DCSBM.
Ideally, the estimated mean matrix $\widehat{\bm{M}}^{(m)}$ should be based on the maximum likelihood. However, for the sake of fair comparison to SVPS, we only consider estimating the mean adjacency matrix through spectral clustering as discussed in Section 3.1, where the spectral clustering method is SCORE or RSC.
Next, we compare SVPS with CBIC and ICL under the previously mentioned three simulation settings, each paired with a corresponding likelihood. In each setting, we experiment on different combinations of $(\rho,r)$ and $K=2,3,\ldots,6$, with spectral clustering method being either SCORE or RSC. In each simulation setup, we generate $100$ independent networks in order to compare different methods in terms of the accuracy rates. As to the choice of thresholds in SVPS, we use $2.05$ in both the Poisson and binomial cases (Simulations 1 and 2), and $2.1$ in the negative binomial case (Simulation 3).
The empirical results of the comparison between SVPS and score based methods are shown in Figure 1. SVPS in general performs much better than the score based methods for the setups under consideration. This result may seem surprising given that the score based methods utilize the likelihood information, while SVPS does not even use the correct variance functions. One possible explanation for score based methods underperforming SVPS is that we fit the models only with spectral methods instead of maximum likelihood estimation with EM algorithms. Again, we choose spectral methods for a fair comparison, since there is no likelihood as a guidance to fit the weighted DCSBM for SVPS.
(a) $\rho=0.04$, $r=4$
(b) $\rho=0.06$, $r=3$
(c) $\rho=0.12$, $r=2$
(d) $\rho=0.04$, $r=4$
(e) $\rho=0.06$, $r=3$
(f) $\rho=0.12$, $r=2$
(g) $\rho=0.04$, $r=4$
(h) $\rho=0.06$, $r=3$
(i) $\rho=0.12$, $r=2$
Figure 1: Comparison between SVPS, CBIC and ICL in terms of accuracy rates. Panels (a)(b)(c) are based on the Poisson case; (d)(e)(f) are based on the binomial case; (g)(h)(i) are based on the negative binomial case.
5.2 Les Misèrable Network
In this subsection, we study the Les Misèrable weighted network compiled in Knuth [1993], also analyzed in the literature of network analysis, see, e.g., Newman and Girvan [2004]; Ball et al. [2011]; Newman and Reinert [2016]. In this network, any two characters (nodes) are connected by a weighted edge representing the number of co-occurrences between the pair in the same chapter of the book. The estimated number of communities by some model based approach in Newman and Reinert [2016] is $6$, which may be reasonable since corresponding major subplots are also identified in their analysis.
With this co-occurrence network, we investigate the difference between the estimated numbers of communities when the network is treated as unweighted and weighted. In the unweighted case, by replacing any positive weight with $1$, we apply CBIC and ICL based on the standard DCSBM, as well as two popular methods for unweighted networks: the Bethe Hessian matrix (BH) [Le and Levina, 2015] and stepwise goodness-of-fit (StGoF) [Jin et al., 2022] to the unweighted adjacency matrix to obtain the estimated number of communities $\widehat{K}$. We use SCORE for node clustering in CBIC, ICL and StGoF. Note that the original StGoF fails to stop in this case, so we choose $\widehat{K}$ as the number of communities which corresponds to the smallest test statistic, as suggested in Jin et al. [2022]. The results of rank selection by these methods are listed in Table 1, which are either $3$ or $4$, much smaller than $6$ as suggested in Newman and Reinert [2016]. We choose $K=3$ and apply RSC and SCORE to the unweighted network, and the clustering results are shown in Figure 2.
Table 1: Estimated $K$ in the unweighted adjacency of Les Misèrable.
BH CBIC ICL StGoF
$\widehat{K}$ 4 3 3 $3$
(a) Unweighted with $K=3$ and RSC
(b) Unweighted with $K=3$ and SCORE
Figure 2: Estimated clusters of nodes in Les Misèrable applying RSC and SCORE to the unweighted adjacency with $K=3$.
Coming back to the original weighted network, we compare the results of rank selection based on SVPS, CBIC and ICL. The implementation of CBIC and ICL are based on the likelihood of Poisson DCSBM, while the clustering and model fitting is based on both SCORE and RSC. For the implementation of SVPS, we first preprocess the weighted network into a regularized adjacency matrix $\bm{A}_{\tau}=\bm{A}+\tau\bm{J}_{n}$, where $\bm{J}_{n}$ is the $n\times n$ all-one matrix and $\tau=0.05,0.1,0.25,0.5$ The threshold for sequential testing in SVPS is 2.05. The comparison results are summarized in Table 2, which shows some consistency among these methods. We also plot in Figure 3 the estimated clusters by applying SCORE and RSC to the weighted network with $K=5,6,7$. It seems that RSC yields more interpretable results than SCORE.
Table 2: Estimated $K$ in Les Misèrable by applying SVPS to the regularized weighted adjacency with different $\tau$ values and CBIC, ICL to the original weighted adjacency with Poisson likelihood; rows correspond to SCORE and RSC as the clustering method.
SVPS ($\tau=0.05$) SVPS ($\tau=0.1$) SVPS ($\tau=0.25$) SVPS ($\tau=0.5$) CBIC ICL
SCORE 7 6 7 6 7 7
RSC 7 7 6 6 7 5
(a) Weighted with $K=5$ and RSC
(b) Weighted with $K=5$ and SCORE
(c) Weighted with $K=6$ and RSC
(d) Weighted with $K=6$ and SCORE
(e) Weighted with $K=7$ and RSC
(f) Weighted with $K=7$ and SCORE
Figure 3: Estimated clusters in Les Misèrable by applying RSC or SCORE to the weighted adjacency with $K=5,6,7$.
6 Discussion
This article studies a method on how to select the number of communities for weighted networks. We first propose a generic weighted DCSBM, in which the mean adjacency matrix is modeled as a DCSBM, and the variance profile is related to the mean adjacency matrix through a variance function, while no likelihood is imposed. Next, we introduce a sequential testing approach for selecting the number of communities. In each step, the mean structure is fitted with some clustering method, and the variance profile matrix can be estimated through the variance function. A key step to construct the testing statistic is variance matrix scaling, which will be helpful in normalizing the adjacency matrix. The test statistic is obtained from the normalized adjacency matrix, which is used to determine the number of communities.
In theory, we show the consistency of our proposed procedure for selecting the number of communities under mild conditions on the weighted DCSBM. In particular, the network is allowed to be sparse. The consistency result is comprised of analysis of both the null and the underfitting cases. The nonsplitting properties studied in Jin et al. [2022] can be extended to the weighted DCSBM, which plays an essential role in our theoretical analysis.
For future work, a number of questions remain), both theoretically and methodologically. In theory, we assume the weighted DCSBM satisfies some balanced conditions for analytical convenience, but this condition is very likely to be further relaxed. Our theoretical results are established for known variance function $\nu(\cdot)$, and it would be interesting to have analogous results for estimated variance function $\hat{\nu}(\cdot)$. It would also be interesting to study the null distribution of $T_{n,m}$ asymptotically, based on which p-values can be obtained in each step. However, we are not able to obtain distributional results for $T_{n,m}$ based on the current tools used in this paper.
A key assumption for our method is that the relationship between the variance profile matrix and the mean adjacency matrix can be characterized by a variance function. It would be interesting to see whether this assumption can be relaxed. Note that the estimation of variance profile is for the purpose of matrix scaling, so that the normalization matrix $\bm{\Psi}$ can be estimated under the null case. It is interesting to investigate whether $\bm{\Psi}$ can be consistently estimated even if the variance profile matrix cannot be consistently estimated. We leave this for future study.
Acknowledgment
Y. Liu and X. Li are partially supported by the NSF via the Career Award DMS-1848575. X. Li would like to thank Tracy Ke, Can Le, Haoran Li, Kaizheng Wang, and Ke Wang for their helpful discussions.
Appendix A Supporting Lemmas
We first cite two results from Landa [2022] about the sensitivity analysis of matrix scaling. Here we only state the results for symmetric matrix scaling with row sums equal to $n$.
Lemma A.1 (Landa [2022]).
Let $\bm{A}$ be an $n\times n$ symmetric matrix with positive entries. Then, there exists a unique positive vector $\bm{x}\in\mathbb{R}^{n}$ satisfying
$x_{i}\left(\sum_{j=1}^{n}A_{ij}x_{j}\right)=1,\quad i=1,\ldots,n.$
Furthermore, denote $a_{\max}=\max_{1\leq i\leq j\leq n}A_{ij}$ and $a_{\min}=\min_{1\leq i\leq j\leq n}A_{ij}$. Then,
$\frac{1}{\sqrt{n}}\frac{\sqrt{a_{\min}}}{a_{\max}}\leq x_{i}\leq\frac{1}{\sqrt{n}}\frac{\sqrt{a_{\max}}}{a_{\min}},\quad i=1,\ldots,n.$
Lemma A.2 (Landa [2022]).
Let $\widetilde{\bm{A}}$ be a symmetric matrix with positive entries. Denote $\tilde{a}_{\max}=\max_{1\leq i\leq j\leq n}\widetilde{A}_{ij}$ and $\tilde{a}_{\min}=\min_{1\leq i\leq j\leq n}\widetilde{A}_{ij}$. Suppose there is a constant $\epsilon\in(0,1)$ and a positive vector $\bm{x}\in\mathbb{R}^{n}$, such that
$\max_{1\leq i\leq n}\left|\sum_{j=1}^{n}x_{i}\widetilde{A}_{ij}x_{j}-1\right|\leq\epsilon.$
Denote $x_{\min}=\min_{1\leq i\leq n}x_{i}$. Then, there exists a positive vector $\tilde{\bm{x}}\in\mathbb{R}^{n}$ such that
$\sum_{j=1}^{n}\tilde{x}_{i}\widetilde{A}_{ij}\tilde{x}_{j}=1,\quad i=1,\ldots,n,$
and
$\max_{1\leq i\leq n}\left|\frac{\tilde{x}_{i}}{x_{i}}-1\right|\leq\frac{\epsilon}{1-\epsilon}+4\epsilon\cdot\frac{n^{3/2}}{x_{\min}^{3}}\cdot\frac{\sqrt{\tilde{a}_{\max}}}{\tilde{a}_{\min}^{2}}.$
As a consequence, if $\bm{x}$ satisfies the equations in Lemma A.1, there holds
$\max_{1\leq i\leq n}\left|\frac{\tilde{x}_{i}}{x_{i}}-1\right|\leq\frac{\epsilon}{1-\epsilon}+4\epsilon\cdot\frac{a_{\max}^{3}}{a_{\min}^{3/2}}\cdot\frac{\tilde{a}_{\max}^{1/2}}{\tilde{a}_{\min}^{2}}.$
The following lemma provides a tail bound of the sum of independent random variables satisfying the Bernstein condition.
Lemma A.3 (Bernstein’s Inequality, Corollary 2.11 of Boucheron et al. [2013]).
Let $X_{1},\ldots,X_{n}$ be independent real-valued random variables. Assume that there exist positive numbers $v$ and $c$ such that $\sum_{i=1}^{n}\operatorname{\mathbb{E}}[X_{i}^{2}]\leq v$ and
$\sum_{i=1}^{n}\operatorname{\mathbb{E}}\left[|X_{i}|^{q}\right]\leq\frac{q!}{2}c^{q-2}v\quad\text{for all intergers }q\geq 3.$
Then, for all $t>0$,
$\operatorname{\mathbb{P}}\left(\sum_{i=1}^{n}\left(X_{i}-\operatorname{\mathbb{E}}\left[X_{i}\right]\right)\geq t\right)\leq\exp\left(-\frac{t^{2}}{2(v+ct)}\right).$
The following lemma given in Latała et al. [2018] bounds the operator norm of heterogeneous random matrices.
Lemma A.4 (Remark 4.12 of Latała et al. [2018]).
Suppose that $\bm{X}$ is a random matrix with independent and centered upper-triangular entries. Define the quantities
[…]
Then, for every $0\leq\epsilon\leq 1$ and $t\geq 0$, we have
$\mathbb{P}\left(\|\bm{X}\|\geq 2(1+\epsilon)\sigma_{\infty}+t\right)\leq n\exp\left(-\frac{\epsilon t^{2}}{C\sigma_{\infty}^{*2}}\right),$
where $C$ is a universal constant.
Lemma A.5.
Under Assumption 1, there holds $\left|\lambda_{K}\left(\bm{M}\right)\right|\geq c_{0}\theta_{\min}^{2}n$.
Proof.
Recall that $\bm{M}=\bm{\Theta}\bm{\Pi}\bm{B}\bm{\Pi}^{\top}\bm{\Theta}$, where $\bm{\Theta}\bm{\Pi}=\left[{\bm{\theta}}_{1},{\bm{\theta}}_{2},\ldots,{\bm{\theta}}_{K}\right]$. Denote $\sigma_{k}\left(\cdot\right)$ as the $k$-th singular value of a matrix. It is easy to observe
[…]
Then by the assumption (13), we have
$|\lambda_{K}(\bm{M})|\geq\left|\lambda_{K}(\bm{B})\right|\cdot\sigma_{K}\left(\bm{\Theta}\bm{\Pi}\right)^{2}\geq c_{0}\theta_{\min}^{2}n.$
Lemma A.6.
Under Assumption 1, the scaling factors satisfy
$\left.1\middle/\left(C\theta_{\min}\sqrt{n}\right)\right.\leq\psi_{i}\leq\left.C\middle/\left(\theta_{\min}\sqrt{n}\right)\right.,\quad i=1,\ldots,n$
for some constant $C$ that only relies on $c_{0}$.
Proof.
In light of Assumption 1, this result is implied by Lemma A.1 directly. ∎
Lemma A.7.
For any fixed $c_{0}>0$ in Assumption 1, there holds
$\left\|\bm{\Psi}^{\frac{1}{2}}(\bm{A}-\bm{M})\bm{\Psi}^{\frac{1}{2}}\right\|\leq 2+o_{P}(1)\quad\text{as~{}}n\rightarrow\infty.$
Proof.
Denote $\bm{E}\coloneqq\bm{\Psi}^{\frac{1}{2}}(\bm{A}-\bm{M})\bm{\Psi}^{\frac{1}{2}}$ whose entries are $E_{ij}=(A_{ij}-M_{ij})\psi_{i}^{\frac{1}{2}}\psi_{j}^{\frac{1}{2}}$. By (9), we have
$\sum_{j=1}^{n}\operatorname{\mathbb{E}}\left[E_{ij}^{2}\right]=1,\quad\forall~{}i=1,\ldots,n.$ (16)
Define
$\widehat{E}_{ij}=E_{ij}\mathbf{1}_{\{|E_{ij}|<1/(\log n)\}}\quad\text{and}\quad\widetilde{E}_{ij}=\widehat{E}_{ij}-\operatorname{\mathbb{E}}\left[\widehat{E}_{ij}\right].$
The resulting matrices are denoted as $\widehat{\bm{E}}$ and $\widetilde{\bm{E}}$, respectively. Our proof strategy is to first bound $\|\widetilde{\bm{E}}\|$, and then $\|\widehat{\bm{E}}\|$, and finally $\|\bm{E}\|$.
Notice that $\widetilde{\bm{E}}$ has independent and mean-centered entries for $i\geq j$. Since
$\left|\widehat{E}_{ij}\right|=\left|E_{ij}\right|\mathbf{1}_{\{|E_{ij}|<1/(\log n)\}}\leq\left|E_{ij}\right|,$
almost surely we have
[…]
This implies that
$\sum_{j=1}^{n}\operatorname{\mathbb{E}}\left[\widetilde{E}_{ij}^{2}\right]\leq 1,\quad\forall~{}i=1,\ldots,n.$
Also, since $\|\widehat{E}_{ij}\|_{\infty}\leq 1/(\log n)$, we have $|\operatorname{\mathbb{E}}[\widehat{E}_{ij}]|\leq 1/(\log n)$, and thus
$\|\widetilde{E}_{ij}\|_{\infty}\leq\|\widehat{E}_{ij}\|_{\infty}+|\operatorname{\mathbb{E}}\widehat{E}_{ij}|\leq 2/(\log n).$
Define the quantities
[…]
Then, apply Corollary A.4 to $\widetilde{\bm{E}}$. For any $t>0$, there holds
$\mathbb{P}\left(\|\widetilde{\bm{E}}\|\geq 2(1+\epsilon)\tilde{\sigma}+t\right)\leq n\exp\left(-\frac{\epsilon t^{2}}{C\tilde{\sigma}_{*}^{2}}\right).$
By letting $\epsilon=t=\log^{-1/4}n$, with the above inequalities, we have
$\|\widetilde{\bm{E}}\|\leq 2+o_{P}(1)\quad\text{as~{}}n\rightarrow\infty.$ (17)
[…]
$\operatorname{\mathbb{E}}\left[E_{ij}\mathbf{1}_{\{|E_{ij}|\geq 1/(\log n)\}}\right]=\operatorname{\mathbb{E}}\left[E_{ij}-\widehat{E}_{ij}\right]=-\operatorname{\mathbb{E}}\left[\widehat{E}_{ij}\right].$
By Cauchy-Schwarz inequality,
$\left(\operatorname{\mathbb{E}}\left[\widehat{E}_{ij}\right]\right)^{2}=\left(\operatorname{\mathbb{E}}\left[E_{ij}\mathbf{1}_{\{|E_{ij}|\geq 1/(\log n)\}}\right]\right)^{2}\leq\operatorname{\mathbb{E}}[E_{ij}^{2}]\operatorname{\mathbb{P}}\left(|E_{ij}|\geq 1/(\log n)\right).$ (18)
By Lemma A.6, there holds
$\operatorname{\mathbb{P}}\left(\left|E_{ij}\right|\geq 1/(\log n)\right)$ $=\operatorname{\mathbb{P}}\left(\left|A_{ij}-M_{ij}\right|\geq\frac{1}{\psi_{i}^{1/2}\psi_{j}^{1/2}\log n}\right)$
$\leq\operatorname{\mathbb{P}}\left(\left|A_{ij}-M_{ij}\right|\geq\frac{\theta_{\min}\sqrt{n}}{C\log n}\right).$
Then, by Assumption 1 and Lemma A.3, we have
$\max_{1\leq i,j\leq n}\operatorname{\mathbb{P}}\left(\left|E_{ij}\right|\geq 1/(\log n)\right)$ $\leq\max_{1\leq i\leq j\leq n}2\exp\left(-\frac{\left(\frac{\theta_{\min}\sqrt{n}}{C\log n}\right)^{2}}{2\left(M_{ij}/c_{0}+R\frac{\theta_{\min}\sqrt{n}}{C\log n}\right)}\right)$
$\leq 2\exp\left(-\frac{C\theta_{\min}n}{\theta_{\min}\log^{2}n+\sqrt{n}\log n}\right)$
$=O(n^{-10}),$ (19)
where the last line is due to (12) in Assumption 1. Then, (16) and (18) imply
$\left\|\operatorname{\mathbb{E}}[\widehat{\bm{E}}]\right\|_{F}^{2}$ $=\sum_{1\leq i,j\leq n}\left(\operatorname{\mathbb{E}}\widehat{E}_{ij}\right)^{2}\leq\sum_{1\leq i,j\leq n}\operatorname{\mathbb{E}}[E_{ij}^{2}]\operatorname{\mathbb{P}}\left(\left|E_{ij}\right|\geq 1/(\log n)\right)=O(n^{-9}).$
Then (17) implies
$\|\widehat{\bm{E}}\|\leq\|\widetilde{\bm{E}}\|+\|\operatorname{\mathbb{E}}[\widehat{\bm{E}}]\|_{F}\leq 2+o_{P}(1).$ (20)
Finally, note that
$\operatorname{\mathbb{P}}\left(\widehat{\bm{E}}\neq\bm{E}\right)$ $=\operatorname{\mathbb{P}}\left(\cup_{1\leq i,j\leq n}\{|E_{ij}|\geq 1/(\log n)\}\right)$
$\leq\sum_{1\leq i,j\leq n}\operatorname{\mathbb{P}}(|E_{ij}|\geq 1/(\log n))$
$\leq n^{2}\max_{1\leq i,j\leq n}\operatorname{\mathbb{P}}(|E_{ij}|\geq 1/(\log n))=O(n^{-8}).$
Then (20) implies $\|\bm{E}\|\leq 2+o_{P}(1)$. ∎
Lemma A.8.
Under Assumption 1, there holds
$\left\|\bm{A}-\bm{M}\right\|=O_{P}(\theta_{\min}\sqrt{n})\quad\text{as~{}}n\rightarrow\infty.$
Proof.
This is a straightforward corollary of Lemmas A.6 and A.7. ∎
Lemma A.9.
Under Assumption 1, we have
$\max_{1\leq i\leq n}\left|\frac{d_{i}}{d_{i}^{*}}-1\right|=o_{P}(1)\quad\text{and}\quad\max_{1\leq k,l\leq K}\left|\frac{\bm{1}_{k}^{\top}\bm{A}\bm{1}_{l}}{\bm{1}_{k}^{\top}\bm{M}\bm{1}_{l}}-1\right|=o_{P}(1).$
Proof.
Note that
$d_{i}-d_{i}^{*}=\sum_{j=1}^{n}(A_{ij}-M_{ij}),$
in which the summands satisfy (15). We also have
$\textrm{var}(d_{i}-d_{i}^{*})=\sum_{j=1}^{n}\nu(M_{ij})\leq\frac{1}{c_{0}}\sum_{j=1}^{n}M_{ij}.$
By Lemma A.3, for any fixed $\epsilon>0$, we have
$\sum_{i=1}^{n}\operatorname{\mathbb{P}}\left(|d_{i}-d_{i}^{*}|\geq\epsilon d_{i}^{*}\right)$ $\leq\sum_{i=1}^{n}2\exp\left(-\frac{\epsilon^{2}{d_{i}^{*}}^{2}}{2\left(\frac{1}{c_{0}}(\sum_{j=1}^{n}M_{ij})+R\epsilon d_{i}^{*}\right)}\right)$
$\leq 2n\exp\left(-C\left(\frac{\epsilon^{2}}{1+\epsilon}\right)\theta_{\min}^{2}n\right)=O(n^{-3}),$
where the last inequality is due to (12) in Assumption 1. Therefore, with probability $1-O(n^{-3})$, we have $\max_{1\leq i\leq n}\left|\frac{d_{i}}{d_{i}^{*}}-1\right|\leq\epsilon$. Given $\epsilon$ can be chosen arbitrarily small, we have $\max_{1\leq i\leq n}\left|\frac{d_{i}}{d_{i}^{*}}-1\right|=o_{P}(1)$.
With a similar argument, we have $\max_{1\leq k,l\leq K}\left|\frac{\bm{1}_{k}^{\top}\bm{A}\bm{1}_{l}}{\bm{1}_{k}^{\top}\bm{M}\bm{1}_{l}}-1\right|=o_{P}(1)$. ∎
For unweighted networks, SCORE has been recently proven in Jin et al. [2022] to enjoy the Nonsplitting Property (NSP) for both the underfitting and the null cases. In other words, the true communities are refinements of the estimated communities with high probability. This property is essential for analyzing sequential testing approaches, since it reduces the number of possibilities of DCSBM fitting. We give a formal definition of this property.
Definition 1 (Nonsplitting Property [Jin et al., 2022]).
Let the ground truth of communities in a network be $\mathcal{N}_{1},\ldots,\mathcal{N}_{K}$. Assume $m\leq K$, and $\widehat{\mathcal{N}}^{(m)}_{1},\ldots,\widehat{\mathcal{N}}^{(m)}_{m}$ are the estimated communities by certain clustering method. We say these estimated communities satisfy the Nonsplitting Property (NSP), if the true communities $\mathcal{N}_{1},\ldots,\mathcal{N}_{K}$ are a refinement of the estimated ones. In other words, for any $k=1,\ldots,K$, there is exactly one $l=1,\ldots,m$, such that $\mathcal{N}_{k}\cap\widehat{\mathcal{N}}^{(m)}_{l}\neq\emptyset$.
Then, we have the following lemma that guarantees the NSP for SCORE under weighted DCSBM.
Lemma A.10.
Under Assumption 1 with any fixed $c_{0}>0$, for any fixed $m\leq K$, SCORE satisfies the NSP with probability $1-O(n^{-3})$.
The proof of this lemma is deferred to Appendix C, which follows the idea in Jin et al. [2022]. A crucial step of the proof is the row-wise bounds of eigenvector perturbations [Jin et al., 2017; Abbe et al., 2020].
Appendix B Proofs of the Main Results
Throughout this section, we use $C$ to represent a constant that only relies on $c_{0}$, whose value may change from line to line.
B.1 Proof of Theorem 4.1
This subsection aims to prove Theorem 4.1, so we assume $m=K$. We omit the superscripts in the estimators indicating the $m$-th step for convenience in notation.
Note that
$T_{n,m}=\left|\lambda_{m+1}\left(\widehat{\bm{\Psi}}^{\frac{1}{2}}\bm{A}\widehat{\bm{\Psi}}^{\frac{1}{2}}\right)\right|$
and
$\widehat{\bm{\Psi}}^{\frac{1}{2}}\bm{A}\widehat{\bm{\Psi}}^{\frac{1}{2}}=\widehat{\bm{\Psi}}^{\frac{1}{2}}\bm{M}\widehat{\bm{\Psi}}^{\frac{1}{2}}+\widehat{\bm{\Psi}}^{\frac{1}{2}}(\bm{A}-\bm{M})\widehat{\bm{\Psi}}^{\frac{1}{2}},$
where $\operatorname{rank}\left(\widehat{\bm{\Psi}}^{\frac{1}{2}}\bm{M}\widehat{\bm{\Psi}}^{\frac{1}{2}}\right)\leq K=m$. Then we have
$T_{n,m}\leq\left\|\widehat{\bm{\Psi}}^{\frac{1}{2}}\left(\bm{A}-\bm{M}\right)\widehat{\bm{\Psi}}^{\frac{1}{2}}\right\|.$
On the other hand,
$\widehat{\bm{\Psi}}^{\frac{1}{2}}\left(\bm{A}-\bm{M}\right)\widehat{\bm{\Psi}}^{\frac{1}{2}}$ $=\left(\widehat{\bm{\Psi}}^{\frac{1}{2}}\bm{\Psi}^{-\frac{1}{2}}\right)\left(\bm{\Psi}^{\frac{1}{2}}(\bm{A}-\bm{M})\bm{\Psi}^{\frac{1}{2}}\right)\left(\bm{\Psi}^{-\frac{1}{2}}\widehat{\bm{\Psi}}^{\frac{1}{2}}\right).$
In order to show that $T_{n,m}\leq 2+o_{p}(1)$, it suffices to show the following inequalities
$\begin{cases}\left\|\bm{\Psi}^{\frac{1}{2}}(\bm{A}-\bm{M})\bm{\Psi}^{\frac{1}{2}}\right\|\leq 2+o_{p}(1),\\ \left\|\widehat{\bm{\Psi}}^{\frac{1}{2}}\bm{\Psi}^{-\frac{1}{2}}-\bm{I}_{n}\right\|\leq o_{p}(1).\end{cases}$
The first inequality follows from Lemma A.7 and the second inequality can be shown by the following lemma.
Lemma B.1.
Under the null case $m=K$, Assumption 1 implies
$\max_{1\leq i\leq n}\left|\frac{\hat{\theta}_{i}}{\theta_{i}}-1\right|=o_{P}(1),\quad\max_{1\leq k,l\leq K}\left|\frac{\widehat{B}_{kl}}{B_{kl}}-1\right|=o_{P}(1),$
$\max_{1\leq i,j\leq n}\left|\frac{\widehat{M}_{ij}}{M_{ij}}-1\right|=o_{P}(1),\quad\max_{1\leq k,l\leq n}\left|\frac{\widehat{V}_{kl}}{V_{kl}}-1\right|=o_{P}(1).$
Furthermore, we have
$\left\|\widehat{\bm{\Psi}}^{\frac{1}{2}}\bm{\Psi}^{-\frac{1}{2}}-\bm{I}_{n}\right\|=o_{P}(1).$
Proof.
In the null case, the NSP given in Theorem A.10 implies strong consistency for the recovery of communities. Without loss of generality, denote $\widehat{\mathcal{N}}_{k}=\mathcal{N}_{k}$ and thereby $\bm{1}_{k}=\hat{\bm{1}}_{k}$ for $k=1,\ldots,K$, with probability $1-O(n^{-3})$. By combining (3), (4), (5) and (6), we get
$\frac{\theta_{i}}{\hat{\theta}_{i}}=\frac{d_{i}^{*}}{d_{i}}\frac{\sum_{j\in\mathcal{N}_{k}}d_{j}}{\sum_{j\in\mathcal{N}_{k}}d_{j}^{*}}\sqrt{\frac{\bm{1}_{k}^{\top}\bm{M}\bm{1}_{k}}{\bm{1}_{k}^{\top}\bm{A}\bm{1}_{k}}},$
and
$\frac{B_{kl}}{\widehat{B}_{kl}}=\frac{\bm{1}_{k}^{\top}\bm{M}\bm{1}_{l}}{\bm{1}_{k}^{\top}\bm{A}\bm{1}_{l}}\sqrt{\frac{\bm{1}_{k}^{\top}\bm{A}\bm{1}_{k}}{\bm{1}_{k}^{\top}\bm{M}\bm{1}_{k}}}\sqrt{\frac{\bm{1}_{l}^{\top}\bm{A}\bm{1}_{l}}{\bm{1}_{l}^{\top}\bm{M}\bm{1}_{l}}}.$
Then by Lemma A.9, we have
$\max_{1\leq i\leq n}\left|\frac{\hat{\theta}_{i}}{\theta_{i}}-1\right|=o_{P}(1),\quad\max_{1\leq k,l\leq K}\left|\frac{\widehat{B}_{kl}}{B_{kl}}-1\right|=o_{P}(1).$
Furthermore, $\max_{1\leq i,j\leq n}\left|\frac{\widehat{M}_{ij}}{M_{ij}}-1\right|=o_{P}(1)$ follows from
$\frac{\widehat{M}_{ij}}{M_{ij}}=\frac{\hat{\theta}_{i}\hat{\theta}_{j}\widehat{B}_{\phi(i)\phi(j)}}{\theta_{i}\theta_{j}B_{\phi(i)\phi(j)}}.$
Then
$\max_{1\leq i,j\leq n}\left|\frac{\widehat{V}_{ij}}{V_{ij}}-1\right|=o_{P}(1)$ (21)
follows from the assumption (14) on the variance-mean function.
On the other hand, notice that (21) implies
$\max_{1\leq i\leq n}\left|\frac{\sum_{j=1}^{n}\widehat{V}_{ij}\psi_{i}\psi_{j}}{\sum_{j=1}^{n}V_{ij}\psi_{i}\psi_{j}}-1\right|=o_{P}(1).$
By (9), we further have
$\max_{1\leq i\leq n}\left|\sum_{j=1}^{n}\widehat{V}_{ij}\psi_{i}\psi_{j}-1\right|=o_{P}(1).$
Denote $V_{\max}=\max_{1\leq i,j\leq n}V_{ij}$ and $V_{\min}=\min_{1\leq i,j\leq n}V_{ij}$. Also, $\widehat{V}_{\max}$ and $\widehat{V}_{\min}$ are similarly defined. Assumption 1 implies
$\left.V_{\max}\middle/V_{\min}\right.=O(1)\quad\text{and}\quad\left.V_{\min}\middle/V_{\max}\right.=O(1).$
Combined with (21), we have
$\left.\widehat{V}_{\max}\middle/V_{\max}\right.=O_{P}(1)\quad\text{and}\quad\left.\widehat{V}_{\min}\middle/V_{\min}\right.=O_{P}(1).$
Plug in the above inequalities to Lemma A.2, we have $\left\|\widehat{\bm{\Psi}}^{\frac{1}{2}}\bm{\Psi}^{-\frac{1}{2}}-\bm{I}_{n}\right\|=o_{P}(1)$. ∎
B.2 Proof of Theorem 4.2
This subsection aims to prove Theorem 4.2 with $m<K$. Again, we omit the superscripts in the estimators.
Note that
$T_{n,m}=\left|\lambda_{m+1}\left(\widehat{\bm{\Psi}}^{\frac{1}{2}}\bm{A}\widehat{\bm{\Psi}}^{\frac{1}{2}}\right)\right|\geq\left|\lambda_{K}\left(\widehat{\bm{\Psi}}^{\frac{1}{2}}\bm{A}\widehat{\bm{\Psi}}^{\frac{1}{2}}\right)\right|\geq|\lambda_{K}(\bm{A})|\hat{\psi}_{\min}\geq(|\lambda_{K}(\bm{M})|-\|\bm{A}-\bm{M}\|)\hat{\psi}_{\min},$ (22)
where $\hat{\psi}_{\min}=\min_{1\leq i\leq n}\hat{\psi}_{i}$. It suffices to find lower bounds for $|\lambda_{K}(\bm{M})|$ and $\hat{\psi}_{\min}$, as well as an upper bound for $\|\bm{A}-\bm{M}\|$.
Lemma B.2.
For the underfitting case $m<K$, denote $\widehat{M}_{\max}=\max_{1\leq i,j\leq n}\widehat{M}_{ij}$ and $\widehat{M}_{\min}=\min_{1\leq i,j\leq n}\widehat{M}_{ij}$. Also, denote $\widehat{V}_{\max}$ and $\widehat{V}_{\min}$ in a similar manner. Then under Assumption 1, there holds
$\max\left(\frac{\widehat{M}_{\max}}{\theta_{\min}^{2}},~{}~{}\frac{\theta_{\min}^{2}}{\widehat{M}_{\min}}\right)=O_{P}(1),$
and
$\max\left(\frac{\widehat{V}_{\max}}{\theta_{\min}^{2}},~{}~{}\frac{\theta_{\min}^{2}}{\widehat{V}_{\min}}\right)=O_{P}(1).$
Furthermore, we have
$\frac{1}{\theta_{\min}\hat{\psi}_{\min}\sqrt{n}}=O_{P}(1).$
Proof.
In the underfitting case $m<K$, the NSP given in Theorem A.10 implies that the true communities $\mathcal{N}_{1},\ldots,\mathcal{N}_{K}$ are refinements of the estimated communities $\widehat{\mathcal{N}}_{1},\ldots,\widehat{\mathcal{N}}_{m}$. For each $1\leq k\leq m$, we assume that the number of true communities contained in $\widehat{\mathcal{N}}_{k}$ is $r_{k}\geq 1$, which implies that $r_{1}+\cdots+r_{m}=K$. Then, we can represent the estimated communities as
$\widehat{\mathcal{N}}_{k}=\mathcal{N}_{h_{k1}}\cup\cdots\cup\mathcal{N}_{h_{kr_{k}}},\quad k=1,\ldots,m.$
Here all indices $h_{kj}$ for $k=1,\ldots,m$ and $j=1,\ldots,r_{k}$ are distinct over $1,\ldots,K$. We can also decompose
$\hat{\bm{1}}_{k}=\bm{1}_{h_{k1}}+\cdots+\bm{1}_{h_{kr_{k}}},\quad k=1,\ldots,m.$
Recall that
$\widehat{M}_{ij}=\frac{(\hat{\bm{1}}_{k})^{\top}\bm{A}\hat{\bm{1}}_{l}}{(\hat{\bm{1}}_{k})^{\top}\bm{A}\bm{1}_{n}\cdot(\hat{\bm{1}}_{l})^{\top}\bm{A}\bm{1}_{n}}d_{i}d_{j}.$
Then by Lemma A.9, NSP and Assumption 1, it is easy to obtain
$\max\left(\frac{\widehat{M}_{\max}}{\theta_{\min}^{2}},~{}\frac{\theta_{\min}^{2}}{\widehat{M}_{\min}}\right)=O_{P}(1).$
Furthermore, by (14) in Assumption 1, we have
$\max\left(\frac{\widehat{V}_{\max}}{\theta_{\min}^{2}},~{}\frac{\theta_{\min}^{2}}{\widehat{V}_{\min}}\right)=O_{P}(1).$
Then by (10) and Lemma A.1, we have
$\frac{1}{\theta_{\min}\hat{\psi}_{\min}\sqrt{n}}=O_{P}(1).$
Proof of Theorem 4.2.
By Assumption 1, we have
$\theta_{\min}\sqrt{n}\geq\log^{3}n\rightarrow\infty.$
Combined with Lemma A.5 and A.8, Theorem 4.2 is proved from (22).
Appendix C Proof of the Nonsplitting Property
The proof of the NSP for SCORE, i.e., Lemma A.10, basically follows the arguments in Jin et al. [2017] and Jin et al. [2022]. In other words, Lemma A.10 is an extension of Theorem 3.2 of Jin et al. [2022] to the case of weighted DCSBM. To carry out this extension, we first need some probabilistic results for subexponential random vectors and matrices.
C.1 Preliminaries
Lemma C.1 (Vector Bernstein-type Inequality, Theorem 2.5 of Bosq [2000]).
[…]
$\sum_{i=1}^{n}\operatorname{\mathbb{E}}\|\bm{X}_{i}\|^{p}\leq\frac{p!}{2}\sigma^{2}R^{p-2},\quad p=2,3,4,\ldots.$
Then, for all $t>0$,
$\operatorname{\mathbb{P}}\left(\left\|\sum_{i=1}^{n}\bm{X}_{i}\right\|\geq t\right)\leq 2\exp\left(-\frac{t^{2}}{2(\sigma^{2}+Rt)}\right).$
The following lemma is an application of the vector Bernstein-type inequality supporting the proof of Lemma C.7.
Lemma C.2.
Let $X_{1},\ldots,X_{n}$ be independent random variables satisfying
$\operatorname{\mathbb{E}}[|X_{i}-\mu_{i}|^{p}]\leq C_{0}\left(\frac{p!}{2}\right)R^{p-2}\mu_{i},$
where $\mu_{\max}\coloneqq\max_{1\leq i\leq n}\mu_{i}\leq C_{1}$. Let $\bm{w}_{1},\ldots,\bm{w}_{n}\in\mathbb{R}^{d}$ be fixed vectors. Then, with probability $1-O(n^{-4})$, we have
$\left\|\sum_{i=1}^{n}(X_{i}-\mu_{i})\bm{w}_{i}\right\|_{2}\leq C\left(\sqrt{\mu_{\max}}\|\bm{W}\|_{F}\sqrt{\log n}+\|\bm{W}\|_{2\rightarrow\infty}(\log n)\right),$ (23)
where $\bm{W}^{\top}=[\bm{w}_{1},\ldots,\bm{w}_{n}]$, $\|\bm{W}\|_{2\rightarrow\infty}=\max_{1\leq i\leq n}\|\bm{w}_{i}\|_{2}$, and $C$ is a constant that only relies on $C_{0}$, $R$, and $C_{1}$.
Proof.
For any $p\geq 2$, we have
$\sum_{i=1}^{n}\operatorname{\mathbb{E}}\|(X_{i}-\mu_{i})\bm{w}_{i}\|_{2}^{p}$ $=\sum_{i=1}^{n}(\operatorname{\mathbb{E}}|X_{i}-\mu_{i}|^{p})\|\bm{w}_{i}\|_{2}^{p}$
$\leq\sum_{i=1}^{n}C_{0}\left(\frac{p!}{2}\right)R^{p-2}\mu_{i}\|\bm{w}_{i}\|_{2}^{p}$
$\leq\sum_{i=1}^{n}C_{0}\left(\frac{p!}{2}\right)R^{p-2}C_{1}\|\bm{w}_{i}\|_{2}^{2}\|\bm{W}\|_{2\rightarrow\infty}^{p-2}$
$=\left(\frac{p!}{2}\right)\left(R\|\bm{W}\|_{2\rightarrow\infty}\right)^{p-2}(C_{0}C_{1}\|\bm{W}\|_{F}^{2}),$
Then, by Lemma C.1, for any $t>0$,
$\operatorname{\mathbb{P}}\left(\left\|\sum_{i=1}^{n}(X_{i}-\lambda_{i})\bm{w}_{i}\right\|_{2}\geq t\right)\leq 2\exp\left(-\frac{t^{2}}{C\mu_{\max}\|\bm{W}\|_{F}^{2}+(R\|\bm{W}\|_{2\rightarrow\infty})t}\right).$
Then we get (23) with probability $1-O(n^{-4})$ for sufficiently large $C$. ∎
The following lemma is the subexponential case of the matrix Bernstein inequality.
Lemma C.3 (Theorem 6.2 of Tropp [2012]).
Consider a finite sequence $\{\bm{X}_{k}\}$ of independent, random, symmetric matrices with dimension $d$. Assume that
$\operatorname{\mathbb{E}}[\bm{X}_{k}]=\bm{0}~{}~{}\text{and}~{}~{}\operatorname{\mathbb{E}}[\bm{X}_{k}^{p}]\preceq\frac{p!}{2}R^{p-2}\bm{A}_{k}^{2},\quad p=2,3,4,\ldots.$
Compute the variance parameter
$\sigma^{2}\coloneqq\left\|\sum_{k}\bm{A}_{k}^{2}\right\|.$
Then, the following chain of inequalities holds for all $t\geq 0$:
$\operatorname{\mathbb{P}}\left(\lambda_{\max}\left(\sum_{k}\bm{X}_{k}\right)\geq t\right)$ $\leq d\exp\left(-\frac{t^{2}}{2(\sigma^{2}+Rt)}\right).$
The following theorem from Abbe et al. [2020] provides perturbation bounds of eigenspaces which is crucial to the proof of Theorem A.10.
Lemma C.4 (Abbe et al. [2020]).
Suppose $\bm{A}\in\mathbb{R}^{n\times n}$ is a symmetric random matrix, and let $\bm{A}^{*}=\operatorname{\mathbb{E}}[\bm{A}]$. Denote the eigenvalues of $\bm{A}$ by $\lambda_{1}\geq\cdots\geq\lambda_{n}$, and their associated eigenvectors by $\{\bm{u}_{j}\}_{j=1}^{n}$. Analogously for $\bm{A}^{*}$, the eigenvalues and eigenvectors are denoted by $\lambda_{1}^{*}\geq\cdots\geq\lambda_{n}^{*}$ and $\{\bm{u}_{j}^{*}\}_{j=1}^{n}$. For convenience, we also define $\lambda_{0}=\lambda_{0}^{*}=\infty$ and $\lambda_{n+1}=\lambda_{n+1}^{*}=-\infty$. Note that we allow eigenvalues to be identical, so some eigenvectors may be defined up to rotations.
Suppose $r$ and $s$ are two integers satisfying $1\leq r\leq n$ and $0\leq s\leq n-r$. Let $\bm{U}=[\bm{u}_{s+1},\ldots,\bm{u}_{s+r}]\in\mathbb{R}^{n\times r}$, $\bm{U}^{*}=[\bm{u}_{s+1}^{*},\ldots,\bm{u}_{s+r}^{*}]\in\mathbb{R}^{n\times r}$ and $\bm{\Lambda}^{*}=\text{diag}(\lambda_{s+1}^{*},\ldots,\lambda_{s+r}^{*})\in\mathbb{R}^{r\times r}$. Define
$\Delta^{*}=(\lambda_{s}^{*}-\lambda_{s+1}^{*})\wedge(\lambda_{s+r}^{*}-\lambda_{s+r+1}^{*})\wedge\min_{1\leq i\leq r}|\lambda_{s+i}^{*}|$
and
$\kappa=\max_{1\leq i\leq r}\left.|\lambda_{s+i}^{*}|\middle/\Delta^{*}\right..$
Suppose for some $\gamma\geq 0$, the following assumptions hold:
  1. A1
    (Incoherence) $\|\bm{A}^{*}\|_{2\rightarrow\infty}\leq\gamma\Delta^{*}$.
  2. A2
[…]
  3. A3
    (Spectral norm concentration) $32\kappa\max\{\gamma,\varphi(\gamma)\}\leq 1$ and for some $\delta_{0}\in(0,1)$,
    $\operatorname{\mathbb{P}}\left(\|\bm{A}-\bm{A}^{*}\|\leq\gamma\Delta^{*}\right)\geq 1-\delta_{0}.$
  4. A4
    (Row concentration) Suppose $\varphi(x)$ is continuous and non-decreasing in $\mathbb{R}_{+}$ with $\varphi(0)=0$, $\varphi(x)/x$ is non-increasing in $\mathbb{R}_{+}$, and $\delta_{1}\in(0,1)$. For any $i\in[n]$ and $\bm{W}\in\mathbb{R}^{n\times r}$,
    $\operatorname{\mathbb{P}}\left(\|(\bm{A}-\bm{A}^{*})_{i\cdot}\bm{W}\|_{2}\leq\Delta^{*}\|\bm{W}\|_{2\rightarrow\infty}\varphi\left(\frac{\|\bm{W}\|_{F}}{\sqrt{n}\|\bm{W}\|_{2\rightarrow\infty}}\right)\right)\geq 1-\frac{\delta_{1}}{n}.$
Under Assumptions (A1)—(A4), with probability at least $1-\delta_{0}-2\delta_{1}$, there exists an orthogonal matrix $\bm{Q}$ such that
$\|\bm{U}\bm{Q}-\bm{A}\bm{U}^{*}(\bm{\Lambda}^{*})^{-1})\|_{2\rightarrow\infty}\lesssim\kappa(\kappa+\varphi(1))(\gamma+\varphi(\gamma))\|\bm{U}^{*}\|_{2\rightarrow\infty}+\gamma\|\bm{A}^{*}\|_{2\rightarrow\infty}\left.\middle/\Delta^{*}\right..$ (24)
Here $\lesssim$ only hides absolute constants.
C.2 Spectral Properties for the Adjacency Matrix
We now introduce some spectral properties of the mean adjacency matrix $\bm{M}$ and the observed weighted adjacency matrix $\bm{A}$, which are useful for proving the NSP for SCORE. We first study the eigenvalues and eigenvectors of $\bm{M}$. The population adjacency matrix $\bm{M}$ is a rank-$K$ matrix, with nonzero eigenvalues $\lambda_{1}^{*},\lambda_{2}^{*},\ldots,\lambda_{K}^{*}$ sorted in descending magnitude, and their corresponding unit-norm eigenvectors $\bm{u}_{1}^{*},\ldots,\bm{u}_{K}^{*}$. Notice that by Perron’s theorem [Horn and Johnson, 2012], $\lambda_{1}^{*}$ is positive with multiplicity 1, and we can choose $\bm{u}_{1}^{*}$ such that all its entries are strictly positive. Denote $\bm{U}^{*}=[\bm{u}_{2}^{*},\ldots,\bm{u}_{K}^{*}]\in\mathbb{R}^{n\times(K-1)}$ and $\left(\bm{U}^{*}_{i\cdot}\right)^{\top}$ as its $i$-th row. Also, denote $\bm{\Lambda}^{*}=\text{diag}(\lambda_{2}^{*},\ldots,\lambda_{K}^{*})\in\mathbb{R}^{(K-1)\times(K-1)}$.
Lemma C.5 (Lemma B.1 of Jin et al. [2017]).
Under Assumption 1, the following statements are true:
  1. (1)
    $|\lambda_{k}^{*}|\asymp\|\bm{\theta}\|_{2}^{2}$, $1\leq k\leq K$.
  2. (2)
    $\lambda_{1}^{*}-|\lambda_{2}^{*}|\asymp\lambda_{1}^{*}$.
Proof.
This result is basically the same as Lemma B.1 of Jin et al. [2017] under Assumption 1. In particular, $\lambda_{1}^{*}-|\lambda_{2}^{*}|\asymp\lambda_{1}^{*}$ can be straightforwardly obtained from Han and Han [2019]. ∎
Also, we have the following properties of the eigenvectors of $\bm{M}$:
Lemma C.6 (Lemma B.2 of Jin et al. [2017]).
Under Assumption 1, the following statements are true:
  1. (1)
    If we choose the sign of $\bm{u}_{1}^{*}$ such that $\sum_{i=1}^{n}u_{1}^{*}(i)>0$, then the entries of $\bm{u}_{1}^{*}$ are positive satisfying $C^{-1}\theta_{i}/\|\bm{\theta}\|_{2}\leq u_{1}^{*}(i)\leq C\theta_{i}/\|\bm{\theta}\|_{2}$, $1\leq i\leq n$.
  2. (2)
    $\|\bm{U}^{*}_{i\cdot}\|_{2}\leq C\sqrt{K}\theta_{i}/\|\bm{\theta}\|_{2}$, $1\leq i\leq n$.
Here $C$ is a constant that only depends on $c_{0}$.
Let’s come back to community detection with SCORE. Let $|\lambda_{1}|\geq|\lambda_{2}|\geq\cdots\geq|\lambda_{K}|$ be the leading $K$ eigenvalues of $\bm{A}$ in magnitude, with corresponding unit-norm eigenvectors $\bm{u}_{1},\ldots,\bm{u}_{K}$. Denote $\bm{U}=[\bm{u}_{2},\ldots,\bm{u}_{K}]\in\mathbb{R}^{n\times(K-1)}$. The proof of the following lemma is similar to that of Lemma 2.1 in Jin et al. [2017], but we need to make some adaptations for weighted DCSBM. We defer the proof to Section C.4.
Lemma C.7.
Under Assumption 1, with probability $1-O(n^{-3})$, the following statements are true:
  1. (1)
    We can select $\bm{u}_{1}$ such that $\|\bm{u}_{1}-\bm{u}_{1}^{*}\|_{\infty}\leq\frac{C(c_{0})}{\sqrt{n\log n}}$.
  2. (2)
    $\|\bm{U}\bm{Q}^{\top}-\bm{U}^{*}\|_{2\rightarrow\infty}\leq\frac{C(c_{0})}{\sqrt{n\log n}}$ for some $\bm{Q}\in\mathcal{O}_{K-1}$.
C.3 Proof of Lemma A.10
With the above lemmas, we can prove Lemma A.10 by following the arguments in Jin et al. [2022]. Here we give a brief outline.
Define $\bm{R}^{(K)}$ as an $n\times(K-1)$ matrix constructed from the eigenvectors $\bm{u}_{1}$, $\bm{u}_{2}$, …, $\bm{u}_{K}$ by taking entrywise ratios between $\bm{u}_{2}$,…, $\bm{u}_{K}$ and $\bm{u}_{1}$, i.e., $R^{(K)}(i,k)=u_{k+1}(i)/u_{1}(i)$ for $1\leq i\leq n$ and $10\leq k\leq K-$ 1. For any $2\leq m\leq K$, let $\bm{R}^{(m)}$ be an $n\times(m-1)$ matrix consists of the first $m-1$ columns of $\bm{R}^{(K)}$. For any candidate number of clusters $m$, SCORE amounts to conducting $k$-means clustering on the rows of $\bm{R}^{(m)}$. Therefore, we need to study the statistical properties of the rows of $\bm{R}^{(m)}$.
Next, we define the population counterparts. Denote $\mathcal{O}_{K-1}$ as the space of $(K-1)\times(K-1)$ orthogonal matrices. For any $\bm{Q}\in\mathcal{O}_{K-1}$ and any $2\leq k\leq K$, let $\bm{u}_{k}^{*}(\bm{Q})$ be the $(k-1)$-th column of $[\bm{u}_{2}^{*},\ldots,\bm{u}_{K}^{*}]\bm{Q}$. Define $\bm{R}^{*(K)}(\bm{Q})\in\mathbb{R}^{n\times(K-1)}$, whose $(k-1)$-th column is the entrywise ratio between $\bm{u}_{k}^{*}(\bm{Q})$ and $\bm{u}_{1}^{*}$. For any $2\leq m\leq K$, let $\bm{R}^{*(m)}(\bm{Q})\in\mathbb{R}^{n\times(m-1)}$ consists of the first $m-1$ columns of $\bm{R}^{*(K)}(\bm{Q})$. The following lemma provides an uniform upper bound between the corresponding rows of $\bm{R}^{(m)}$ and $\bm{R}^{*(m)}(\bm{Q})$.
Lemma C.8 (Row-wise deviation bound).
For $2\leq m\leq K$, denote $\left(\bm{r}_{i}^{(m)}\right)^{\top}$ and $\left(\bm{r}_{i}^{*(m)}(\bm{Q})\right)^{\top}$ as the $i$-th row of $\bm{R}^{(m)}$ and $\bm{R}^{*(m)}(\bm{Q})$, respectively. Under Assumption 1, with probability $1-O(n^{-3})$, there exists a $(K-1)\times(K-1)$ orthogonal matrix $\bm{Q}$, such that
$\|\bm{r}_{i}^{(m)}-\bm{r}_{i}^{*(m)}(\bm{Q})\|_{2}\leq\|\bm{r}_{i}^{(K)}-\bm{r}_{i}^{*(K)}(\bm{Q})\|_{2}\leq\frac{C(c_{0})}{\sqrt{\log n}},$ (25)
for each $2\leq m\leq K$ and $1\leq i\leq n$.
This lemma is a straightforward corollary of Lemmas C.6 and C.7 by following the proof of Lemma 4.1 in Jin et al. [2022], so we omit the details.
Combining Lemma C.8 with Lemma 4.2, Lemma 4.3 and Theorem 4.1 in Jin et al. [2022], Lemma A.10 is proved by exactly the same arguments as in Jin et al. [2022].
C.4 Proof of Lemma C.7
Proof.
Divide $\lambda_{1}^{*},\ldots,\lambda_{K}^{*}$ into three groups: (1) $\lambda_{1}^{*}$, (2) positive values in $\lambda_{2}^{*},\ldots,\lambda_{K}^{*}$, and (3) negative values in $\lambda_{2}^{*},\ldots,\lambda_{K}^{*}$. We shall apply Lemma C.4 to all the three groups. For succinctness, we only show in detail applying the theorem to group (2), while the proof for the other two groups is similar and thus omitted.
Denote $K_{1}$ as the number of eigenvalues in group (2). Define $\bm{\Lambda}_{1}^{*}$ as the diagonal matrix consisting of eigenvalues in group (2), and $\bm{U}_{1}^{*}$ as the matrix whose columns are the associated eigenvectors. Define the two matrices’ empirical counterparts as $\bm{\Lambda}_{1}$ and $\bm{U}_{1}$. To show the second statement in the lemma, we aim to first show with high probability that
  1. (i)
    $\|\bm{U}_{1}\bm{Q}^{\top}-\bm{A}\bm{U}_{1}^{*}(\bm{\Lambda}_{1}^{*})^{-1}\|_{2\rightarrow\infty}\leq\frac{C(c_{0})}{\sqrt{n\log n}}$ for some $\bm{Q}\in\mathcal{O}_{K_{1}}$.
  2. (ii)
    $\|\bm{U}_{1}^{*}-\bm{A}\bm{U}_{1}^{*}(\bm{\Lambda}_{1}^{*})^{-1}\|_{2\rightarrow\infty}\leq\frac{C(c_{0})}{\sqrt{n\log n}}$.
Proof of (i)
To apply Lemma C.4, we need to determine $\gamma$ and $\varphi:\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$, and then verify Assumption (A1)—(A4) as required by the theorem. Note that by Lemma C.5, we have
$\Delta^{*}=\min\left\{\lambda_{1}^{*}-\lambda_{2}^{*},|\lambda_{K}^{*}|\right\}\geq C(c_{0})\theta_{\min}^{2}n,$
and $\kappa\leq C(c_{0})$. We choose an appropriately large $C_{1}(c_{0})$ and let
$\gamma=\frac{C_{1}(c_{0})}{\sqrt{\log n}}.$
  1. A1
    This is trivial since
    $\|\bm{M}\|_{2\rightarrow\infty}\leq c_{0}^{-2}\theta_{\min}^{2}\sqrt{n}\leq\Delta^{*}\gamma,$
    when $n$ is sufficiently large.
  2. A2
    This obviously holds.
  3. A3
    In the proof of A4, we show that $\varphi(\gamma)\leq C(c_{0})(\log n)^{-3/2}$, hence, $32\kappa\max\{\gamma,\varphi(\gamma)\}\leq 1$ when $n$ is sufficiently large.
    Next, we show the spectral norm perturbation bound by applying the subexponential case of matrix Bernstein inequality (Theorem 6.2 of Tropp [2012], restated as Lemma C.3 in this paper). Let $\bm{X}^{ij}=(A_{ij}-M_{ij})(\bm{E}^{ij}+\bm{E}^{ji})$ for $i<j$, and $\bm{X}^{ii}=(A_{ii}-M_{ii})\bm{E}^{ii}$ for $i=1,\ldots,n$, where $\bm{E}^{ij}$ is a $n\times n$ matrix with 1 on the $(i,j)$-th entry and 0 elsewhere. Obviously, $\operatorname{\mathbb{E}}[\bm{X}^{ij}]=\bm{0}$. Notice that
    $\operatorname{\mathbb{E}}[(\bm{X}^{ij})^{p}]=\begin{cases}\operatorname{\mathbb{E}}[(A_{ij}-M_{ij})^{p}](\bm{E}^{ij}+\bm{E}^{ji})\quad\text{when $p$ is odd};\\ \operatorname{\mathbb{E}}[(A_{ij}-M_{ij})^{p}](\bm{E}^{ij}+\bm{E}^{ji})\quad\text{when $p$ is even}.\end{cases}$
    Also, note that $-(\bm{E}^{ii}+\bm{E}^{jj})\preceq\bm{E}^{ij}+\bm{E}^{ji}\preceq\bm{E}^{ii}+\bm{E}^{jj}$. Then, by (15) in Assumption 1, for any integer $p\geq 2$, we have
    $\left|\operatorname{\mathbb{E}}[(A_{ij}-M_{ij})^{p}]\right|\leq\operatorname{\mathbb{E}}[|A_{ij}-M_{ij}|^{p}]\leq C^{\prime}\left(\frac{p!}{2}\right)R^{p-2}M_{ij},$
    where $C^{\prime}$ and $R$ only depend on $c_{0}$. Then,
    $\operatorname{\mathbb{E}}[(\bm{X}^{ij})^{p}]\preceq C^{\prime}\frac{p!}{2}R^{p-2}M_{ij}(\bm{E}^{ii}+\bm{E}^{jj}),\quad p=2,3,4,\ldots.$
    Thus, the conditions of Lemma C.3 are verified. Notice that $\sum_{i\leq j}\bm{X}^{ij}=\bm{A}-\bm{M}$. Denote
[…]
    where $\lesssim$ only hides a constant depending on $c_{0}$. Then, for all $t\geq 0$,
    $\operatorname{\mathbb{P}}\left(\|\bm{A}-\bm{M}\|\geq t\right)\leq n\exp\left(-\frac{t^{2}}{2(\sigma^{2}+Rt)}\right)\leq n\exp\left(-\frac{1}{4}\left(\frac{t^{2}}{\sigma^{2}}\wedge\frac{t}{R}\right)\right).$
    By (12) in Assumption 1, when $n$ is sufficiently large, with probability $1-O(n^{-3})$, for a sufficiently large $C(c_{0})$,
    $\|\bm{A}-\bm{M}\|\leq C(c_{0})\theta_{\min}\sqrt{n\log n}.$
  4. A4
    By Lemma C.2, for any $1\leq i\leq n$ and $\bm{W}\in\mathbb{R}^{n\times r}$, with probability $1-O(n^{-4})$, there holds
    $\|(\bm{A}-\bm{M})_{i\cdot}\bm{W}\|_{2}\leq C_{2}(c_{0})\max\left(\theta_{\min}\|\bm{W}\|_{F}\sqrt{\log n},\|\bm{W}\|_{2\rightarrow\infty}(\log n)\right),$ (26)
    for a sufficiently large $C_{2}(c_{0})$. Since $\Delta^{*}\geq C(c_{0})\theta_{\min}^{2}n$, we have
    $\|(\bm{A}-\bm{M})_{i\cdot}\bm{W}\|_{2}$ $\leq C_{2}(c_{0})\Delta^{*}\|\bm{W}\|_{2\rightarrow\infty}\max\left(\frac{\sqrt{n\log n}}{\theta_{\min}n}\frac{\|\bm{W}\|_{F}}{\sqrt{n}\|\bm{W}\|_{2\rightarrow\infty}},\frac{\log n}{\theta_{\min}^{2}n}\right).$
    Now we follow similar arguments as in the proof of Lemma 2.1 of Jin et al. [2017]. Define the quantities $t_{1}=C_{2}(c_{0})\left(\theta_{\min}n\right)^{-1}\sqrt{n\log n}$ and $t_{2}=C_{2}(c_{0})\left(\theta_{\min}^{2}n\right)^{-1}\log n$ Define the function
    $\widetilde{\varphi}(x)=\max\left(t_{1}x,t_{2}\right).$
    Then, we have for any $1\leq i\leq n$, with probability $1-O(n^{-4})$,
    $\|(\bm{A}-\bm{M})_{i\cdot}\bm{W}\|_{2}\leq\Delta^{*}\|\bm{W}\|_{2\rightarrow\infty}\widetilde{\varphi}\left(\frac{\|\bm{W}\|_{F}}{\sqrt{n}\|\bm{W}\|_{2\rightarrow\infty}}\right).$ (27)
    First, notice that $(\sqrt{n}\|\bm{W}\|_{2\rightarrow\infty})^{-1}\|\bm{W}\|_{F}\in\left[n^{-1/2},1\right]$. Next, observe that by (12), when $n$ is sufficiently large, we have $t_{2}/t_{1}=(\theta_{\min}\sqrt{n})^{-1}\sqrt{\log n}>n^{-1/2}$. Therefore, when $x\in\left[n^{-1/2},t_{2}/t_{1}\right]$, $t_{1}x\leq t_{2}$, i.e., $\widetilde{\varphi}(x)=t_{2}$; when $x\in\left(t_{2}/t_{1},1\right]$, $t_{1}x>t_{2}$, i.e., $\widetilde{\varphi}(x)=t_{1}x$. As a result, we can construct $\varphi(\cdot)$ as:
    $\varphi(x)=\begin{cases}\sqrt{n}t_{2}x\quad&\text{for}~{}0\leq x\leq n^{-1/2};\\ t_{2}\quad&\text{for}~{}n^{-1/2}<x\leq t_{2}/t_{1};\\ t_{1}x\quad&\text{for}~{}t_{2}/t_{1}<x\leq 1;\\ t_{1}\quad&\text{for}~{}x>1.\end{cases}$
    Obviously, $\varphi(x)$ is continuous and non-decreasing in $\mathbb{R}_{+}$, with $\varphi(0)=0$ and $\varphi(x)/x$ being non-increasing in $\mathbb{R}_{+}$. By (27) and $0\leq\widetilde{\varphi}(x)\leq\varphi(x)$, we have with probability $1-O(n^{-4})$,
    $\|(\bm{A}-\bm{M})_{i\cdot}\bm{W}\|_{2}\leq\Delta^{*}\|\bm{W}\|_{2\rightarrow\infty}\varphi\left(\frac{\|\bm{W}\|_{F}}{\sqrt{n}\|\bm{W}\|_{2\rightarrow\infty}}\right).$
    By the definition of $t_{1}$ and (12), we have $t_{1}\leq(\log n)^{-5/2}$. Furthermore, since $\varphi(x)\leq t_{1}$, we have $\varphi(\gamma)\leq(\log n)^{-5/2}$.
After verifying (A1)—(A4), we obtain the bound (24). Note that based on the definition of $\varphi(x)$, we have $\kappa(\kappa+\varphi(1))\leq C(c_{0})$ and $\gamma+\varphi(\gamma)\leq\frac{C(c_{0})}{\sqrt{\log n}}$. Also, by Lemma C.6, $\|\bm{U}^{*}_{1}\|_{2\rightarrow\infty}\leq\frac{C(c_{0})}{\sqrt{n}}$. Since $\|\bm{M}\|_{2\rightarrow\infty}\leq c_{0}^{-2}\theta_{\min}^{2}\sqrt{n}$ and $\Delta^{*}\geq C(c_{0})\theta_{\min}^{2}n$, we have $\|\bm{M}\|_{2\rightarrow\infty}/\Delta^{*}\leq\frac{C(c_{0})}{\sqrt{n}}$. Finally, we obtain that with probability $1-O(n^{-3})$,
$\left\|\bm{U}_{1}\bm{Q}-\bm{A}\bm{U}^{*}_{1}(\bm{\Lambda}^{*}_{1})^{-1}\right\|_{2\rightarrow\infty}$ $\lesssim\kappa(\kappa+\varphi(1))(\gamma+\varphi(\gamma))\left\|\bm{U}^{*}_{1}\right\|_{2\rightarrow\infty}+\gamma\left.\|\bm{M}\|_{2\rightarrow\infty}\middle/\Delta^{*}\right.$
$\lesssim\frac{1}{\sqrt{n\log n}},$
where $\lesssim$ only hides a constant depending on $c_{0}$ in Assumption 1.
Proof of (ii)
Based on the fact $\bm{U}^{*}_{1}=\bm{M}\bm{U}^{*}_{1}(\bm{\Lambda}^{*}_{1})^{-1}$, we have
$\left\|\bm{U}^{*}_{1}-\bm{A}\bm{U}^{*}_{1}\left(\bm{\Lambda}^{*}_{1}\right)^{-1}\right\|_{2\rightarrow\infty}=\left\|(\bm{M}-\bm{A})\bm{U}^{*}_{1}\left(\bm{\Lambda}^{*}_{1}\right)^{-1}\right\|_{2\rightarrow\infty}.$
By Lemma C.6 and assumption (12), applying (26) with $\bm{W}=\bm{U}^{*}_{1}$ yields that with probability $1-O\left(n^{-3}\right)$,
$\left\|(\bm{A}-\bm{M})\bm{U}^{*}_{1}(\bm{\Lambda}^{*}_{1})^{-1}\right\|_{2\rightarrow\infty}$ $\leq\left(\underset{1\leq i\leq n}{\max}\left\|(\bm{A}-\bm{M})_{i\cdot}\bm{U}^{*}_{1}\right\|_{2}\right)\cdot\left\|\left(\bm{\Lambda}^{*}_{1}\right)^{-1}\right\|$
$\leq C(c_{0})\max\left(\theta_{\min}\left\|\bm{U}^{*}_{1}\right\|_{F}\sqrt{\log n},\text{~{}}\left\|\bm{U}^{*}_{1}\right\|_{2\rightarrow\infty}(\log n)\right)\cdot\left(\theta_{\min}^{2}n\right)^{-1}$
$\leq C(c_{0})\max\left(\left(\theta_{\min}\sqrt{n}\right)^{-1}\left\|\bm{U}^{*}_{1}\right\|_{F}\sqrt{\frac{\log n}{n}},\text{~{}}\left(\theta_{\min}\sqrt{n}\right)^{-2}\left\|\bm{U}^{*}_{1}\right\|_{2\rightarrow\infty}(\log n)\right)$
$\lesssim\frac{1}{\sqrt{n\log^{5}n}},$
where $\lesssim$ only hides a constant depending on $c_{0}$ in Assumption 1.
With (i) and (ii), we have
$\left\|\bm{U}_{1}\bm{Q}^{\top}-\bm{U}_{1}^{*}\right\|_{2\rightarrow\infty}\lesssim\frac{1}{\sqrt{n\log n}}\quad\text{for some~{}}\bm{Q}\in\mathcal{O}_{K_{1}}.$
For eigenvalues in group (3), we similarly have
$\left\|\bm{U}_{2}\bm{Q}^{\top}-\bm{U}_{2}^{*}\right\|_{2\rightarrow\infty}\lesssim\frac{1}{\sqrt{n\log n}}\quad\text{for some~{}}\bm{Q}\in\mathcal{O}_{K_{2}}.$
Combining these results yields the second statement in this lemma: With probability $1-O(n^{-3})$, we have
$\left\|\bm{U}\bm{Q}^{\top}-\bm{U}^{*}\right\|_{2\rightarrow\infty}\lesssim\frac{1}{\sqrt{n\log n}}\quad\text{for some }\bm{Q}\in\mathcal{O}_{K-1}.$
To show the first statement, we apply Lemma C.4 to group (1) with $s=0$ and $r=1$. Note that by Lemma C.5, we have
$\Delta^{*}=\min\{\lambda_{1}^{*},\lambda_{1}^{*}-\lambda_{2}^{*}\}\geq C(c_{0})\theta_{\min}^{2}n$
and $\kappa\leq C(c_{0})$. Following similar procedures, we can select $\bm{u}_{1}$ such that
  1. (i)
    $\left\|\bm{u}_{1}-\left.\bm{A}\bm{u}_{1}^{*}\middle/\lambda_{1}^{*}\right.\right\|_{\infty}\leq\frac{C(c_{0})}{\sqrt{n\log n}}$.
  2. (ii)
    $\left\|\bm{u}_{1}^{*}-\left.\bm{A}\bm{u}_{1}^{*}\middle/\lambda_{1}^{*}\right.\right\|_{\infty}\leq\frac{C(c_{0})}{\sqrt{n\log n}}$.
Therefore, with probability $1-O(n^{-3})$, we have
$\left\|\bm{u}_{1}-\bm{u}_{1}^{*}\right\|_{\infty}\leq\frac{C(c_{0})}{\sqrt{n\log n}},$
which proves the first statement. ∎